OpenStudy (anonymous):

x^3-144x (less than or equal to) 0 anyone?

OpenStudy (anonymous):

x<=12

OpenStudy (anonymous):

is that the final answer?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so how did you solve it

OpenStudy (anonymous):

x^3-144x <= 0 x(x^2-144) <=0 x^2-144 <=0 x^2 <=144 x <= sqrt(144)=12

OpenStudy (anonymous):

i'm not sure but i think the solution set should include the infinity sign in it..

OpenStudy (anonymous):

it can be written as \[(\infty,12] \]

OpenStudy (anonymous):

ok thanks!!!